If 2/5 of a number is 30, what is 1/3 of twice the number?: If 2/5 of a number is 30, what is 1/3 of twice the number?
Answer by nerdybill(1129)

If 2/5 of a number is 30, what is 1/3 of twice the number?
.
Let x = "the number"
.
From: "2/5 of a number is 30" we get our equation:
(2/5)x = 30
2x = 150
x = 75
.
Since we now know "the number" we can answer the question of:
"what is 1/3 of twice the number?"
(1/3)(2x)
= (1/3)(2(75))
= (1/3)(150)
= 50 (this is your answer)